# Computational techniques for sheaf cohomology of locally profinite sets

### v1

@@ Abstract.
We compute the sheaf cohomology with constant $\mathbb{Z}_{2}$ coefficients of a concrete class of locally profinite sets of independent interest. We introduce $k$-Fubini partitions to aid in constructions, which witness a failure of a Fubini theorem analog for these spaces. It is also shown that questions of intermediate cohomology degrees can be reduced to questions about top cohomology degrees by exhibiting nontrivial top cocycles as pointwise limits of coboundaries.
@@ 1. Introduction
The paper contains three main parts, corresponding to the three main tools developed. After expositing the spaces under study in section 2, in section 3 we analyze their sheaf cohomology directly by exhibiting nontrivial cocycles via a nonconstructive argument. Here our arguments are within ZFC, but our conclusions are strengthened by cardinal arithmetic bounds. In section 4, we define $k$-Fubini partitions of the spaces, which witness a failure of a certain analog of Fubini’s theorem. We use these partitions to show both vanishing and nonvanishing of certain cohomology groups; we also prove a converse to one of Aoki’s main results. Lastly, in section 5 we characterize nonvanishing cohomology purely in terms of nonvanishing top cohomology. A better understanding of these top cohomology groups is thus sufficient to resolve all remaining cases. All our arguments may be construed as taking place in ZF, with choice only needed in a few places, such as in the existence of an injection $\aleph_{1}\hookrightarrow 2^{\aleph_{0}}$.
@@ 2. Definitions and overview of results
In section 4, we describe techniques for eliminating the cardinal arithmetic assumptions of Theorem A in certain situations. The main tool is a partition of the product $\prod_{i\leq n}\kappa_{i}$ into $n+1$ pieces, each of which is “thin” in one of the coordinate directions. The situation resembles a failure of Fubini’s theorem, so we call such partitions of the spaces “$k$-Fubini”.
@@
Let $\vec{\kappa}$ be a length-$(n+1)$ tuple of cardinals which admits a $k$-Fubini partition. Then:
@@
We determine exactly when $k$-Fubini partitions exist, in the process deriving a converse to one of Aoki’s results:
@@ 3. Proof of Theorem A
    If $A\subseteq A^{\prime}$, then $\mathcal{D}_{A^{\prime}}\subseteq\mathcal{D}_{A}$ is open dense, so that the map $f\mapsto f\restriction\mathcal{D}_{A}$ defines a dense embedding of topological rings $\mathcal{C}_{A}\hookrightarrow\mathcal{C}_{A^{\prime}}$. In light of this and related observations, we will frequently abuse notation when dealing with functions defined on different domains; for instance, given functions $f_{A}\in\mathcal{D}_{A}$ and $f_{B}\in\mathcal{D}_{B}$, we write $f_{A}+f_{B}$ for the function defined on the intersection $\mathcal{D}_{A}\cap\mathcal{D}_{B}=\mathcal{D}_{A\cup B}$.
@@
    This is a disjoint union of $R$ and finitely many spaces of the form $\prod_{i\in A}J_{i}$, where $A\subseteq n+1$ is of size <n+1<n+1. Since $J_{i}$ is a cofinite subset of $\alpha(\kappa_{i})$ containing $\infty$, any space of this form is homeomorphic to $\mathcal{C}_{\emptyset}(\vec{\kappa}\restriction_{A})$, which by the inductive hypothesis has cardinality $\kappa_{\max A}$.
    We have thus described a bijection
    $\mathcal{C}_{\emptyset}(\vec{\kappa})\xrightarrow{\sim}\prod_{i\leq n}[\kappa_{i}]^{<\aleph_{0}}\times\{0,1\}\times\prod_{A\in[n+1]^{<n+1}}\kappa_{\max A},$
@@
We are now ready to prove Theorem A. We will benefit from the following notation: for a fixed tuple $\vec{\kappa}$, write
@@
That is, if $\vec{y}\in X(\kappa_{n-k+1},\ldots,\kappa_{n})$ satisfies $y_{i}=\infty$ iff $i=j$, then $g_{j}(\vec{x},-)$ must be constant in a neighborhood of $\vec{y}$. Thus there is $\beta(\vec{x},j)\in\kappa_{j}$ such that $g_{j}(\vec{x},\vec{y})$ is independent of $y_{j}$ whenever $y_{j}>\beta(\vec{x},j)$. Taking
@@
we find that $\sum_{j=n-k+1}^{n}g_{j}(\vec{x},\vec{y})$ is independent of $\vec{y}$ whenever $\vec{y}>\vec{\beta}(\vec{x})$ coordinatewise.
The above describes a map $\vec{x}\mapsto\vec{\beta}(\vec{x}):\mathcal{P}\to\mathcal{Q}$; so again by our assumption the image of this map must be bounded in $\mathcal{Q}$. Taking $\vec{\beta}$ to be a common upper bound in $\mathcal{Q}$ of all the $\vec{\beta}(\vec{x})$, we find that $\sum_{j=n-k+1}^{n}g_{j}(\vec{x},\vec{y})$ is constant for $\vec{y}>\vec{\beta}$.
@@
4. Constructions via Fubini partitions
As mentioned above, our second main computational tool is the notion of a $k$-Fubini partition of $X(\vec{\kappa})^{-}$. To define this, we begin with a more careful dissection of the topology of $X(\vec{\kappa})$. We make the following notation:
@@ 4. Constructions via Fubini partitions
$\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}}):=\{\vec{y}\in X(\vec{\kappa}):\vec{y}\upharpoonright_{A}=\vec{x}\upharpoonright_{A}\text{ and }y_{j}=\infty\text{ or }y_{j}\not\in\mathcal{E}_{j}\text{ for all }j\not\in A\}.$
@@
Let $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n}\rangle$ be a tuple of cardinals, and fix $1\leq k\leq n$. We call an $(n+1)$-element partition $\langle Y_{i}:i\leq n\rangle$ of $X(\vec{\kappa})^{-}$ $k$-Fubini if, for any $A\in[n+1]^{\geq k+1}$, the following criteria hold:
@@
    However, it is the more verbose condition (2) which will be most useful to us.
  • It is immediate from the definition that a $k$-Fubini partition is $\ell$-Fubini whenever $k\leq\ell$; moreover, if $\vec{\kappa}\neq\emptyset$ then (1)(-1)-Fubini partitions never exist, and any partition into clopen sets is $n$-Fubini. Consequently, for every nonempty tuple there is a $k_{0}$ such that a $k$-Fubini partition exists if and only if $k_{0}\leq k$.
  • Lastly, we observe that an (n1)(n-1)-Fubini partition is the same as a partition of the product $\prod_{i\leq n}\kappa_{i}$ into $n+1$ many sets $Y_{i}$ such that $Y_{i}$ is “finite in ii”– in the sense that for any $\vec{x}\in\prod_{j\neq i}\kappa_{i}$, there are only finitely many $\vec{y}\in Y_{i}$ extending $\vec{x}$. This is the sense in which the partition witnesses a failure of Fubini’s theorem for $X(\vec{\kappa})$.
We defer discussion of the (non-)existence of Fubini partitions to the end of the section, focusing instead on their uses.
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})$. Then for any $A\in[n+1]^{\geq k+1}$ and any $i\not\in A$,
@@
Fix $A\in[n+1]^{\geq k+1}$ and $i\not\in A$. For any $B\supseteq A$ we have by Lemma 4.3 that $F_{B}\subseteq\cup_{i\in B}Y_{i}$. In particular, if $i\not\in B$ then $F_{B}$ and $Y_{i}$ are disjoint; that is,
@@
As a first example of this, we construct a trivializer of a $k$-cocycle using a $k$-Fubini partition:
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})$, and suppose $\vec{f}\in\prod_{A\in[n+1]^{k+2}}\mathcal{C}_{A}(\vec{\kappa})$ is a tuple of continuous functions. For any $B\in[n+1]^{k+1}$, any $i\leq n$, and any $\vec{x}\in Y_{i}$, define
@@
If there is a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ then $H^{\ell}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ for all $\ell\geq k$.
@@
Figure 2. Given a 2-cocycle $\vec{f}=\langle f_{012},f_{013},f_{023},f_{123}\rangle$ and a $k$-Fubini partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle g_{B}:B\in[4]^{2}\rangle$ defines a trivializer of $\vec{f}$. (For brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
We next use a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ to construct an embedding of the group $H^{k}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})$ into $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ when $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise. A priori, one might try to construct an embedding by treating large enough coordinates as infinity and mapping unrestricted values to zero, as follows:
@@
where $\vec{x}^{\infty}$ is obtained from $\vec{x}$ by setting all $x_{i}>=\lambda_{i}$ equal to $\infty$.
@@
Figure 3 shows an example of the naive extension of a cochain from $X(\vec{\lambda})^{-}$ to $X(\vec{\kappa})^{-}$ for $\lambda\leq\kappa$.
@@
Thus in order to construct an embedding, we will modify the cochain $\vec{f}^{\infty}$ to a cocycle. In fact we will show that, given the existence of a $k$-Fubini partition, any $k$-cochain can be modified to a $k$-cocycle without affecting its values at any limit points.
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-Fubini partition of $X(\vec{\kappa})^{-}$, and let $\vec{f}\in H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ be a $k$-cochain. Define the modification of $\vec{f}$ as follows: for any $A\in[n+1]^{k+1}$ and any $\vec{x}\in\mathcal{D}_{A}\cap Y_{i}$, define
@@
Lastly, to see that modification preserves values at limit points, let $\vec{x}\in\mathcal{D}_{A}$ be a limit point, i.e., $\vec{x}\in F_{B}$ for some $A\subseteq B\subsetneq n+1$. Since $|B|\geq k+1$, condition (2) of the $k$-Fubini partition implies the existence of a neighborhood of $\vec{x}$ in which every point lies in $Y_{i}$ for some $i\in A$. Thus, in this neighborhood $f^{\operatorname{mod}}_{A}\equiv f_{A}.$  ∎
@@
Figure 4. Given a 1-cochain $\vec{f}=\langle f_{01},f_{02},f_{03},f_{12},f_{13},f_{23}\rangle$ and a $k$-Fubini partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle f^{\operatorname{mod}}_{A}:A\in[4]^{2}\rangle$ modifies $\vec{f}$ to a cocycle without affecting its values at any limit points. (Again, for brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
@@
Suppose $X(\vec{\kappa})^{-}$ admits a $k$-Fubini partition. Then for any $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise, there is an embedding
@@
It now remains to determine when Fubini partitions exist. As we saw above, the interesting cases are $0\leq k<n$. We first consider when k<n1k<n-1.
@@
Let $\vec{\kappa}$ be a weakly increasing tuple of cardinals, $0\leq k<n-1$. Then a $k$-Fubini partition of $X(\vec{\kappa})^{-}$ exists if and only if $\kappa_{n-k}=\aleph_{0}$, that is, the first nk+1n-k+1 entries of $\vec{\kappa}$ are all $\aleph_{0}$.
For the forward direction, suppose $\vec{\kappa}$ admits a $k$-Fubini partition. Then we simultaneously have that (i) $\vec{\kappa}$ also admits a $(k+1)$-Fubini partition (Remark 4.3), and that (ii) $H^{k+1}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ (Corollary 4.6). The only way to reconcile these with Proposition 4.11 is if no $\vec{\lambda}\leq\vec{\kappa}$ has nonzero $(k+1)$-cohomology. But now consider the tuple
@@
Then $\mathcal{N}(\vec{x},\vec{\mathcal{E}})$ is a neighborhood of $\vec{x}$, all of whose points have the same minimum. If $\vec{x}\in Y_{i}$ then this neighborhood lies within $Y_{i}$ and if $\vec{x}\not\in Y_{i}$ it lies outside $Y_{i}$. So $Y_{i}$ is clopen in $\mathcal{D}_{A}$. A similar approach shows (2): fix $\vec{x}\in F_{A}$ and for $j\not\in A$ set $\mathcal{E}_{j}=\{0,1,\ldots,x_{i_{0}}\}$. If $\vec{y}\in\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$, then for every $j\not\in A$ we have $y_{j}>x_{i_{0}}=y_{i_{0}}$; so $\min\vec{y}\neq y_{j}$ and thus $\vec{y}\not\in Y_{j}$. We conclude that the minimum of $\vec{y}$ is among the $y_{i}$ for $i\in A$, i.e.,
@@
The remaining case is k=n1k=n-1. Here the backward direction of Proposition 4.12 still holds: if $\kappa_{1}=\aleph_{0}$ then the above defines an (n1)(n-1)-Fubini partition, thus trivializing the top cohomology of $X(\vec{\kappa})^{-}$. On the other hand, Aoki showed that if $\kappa_{i}\geq\aleph_{i}$ for all $i$, then $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$. Our converse to Aoki’s result comes from the requisite strengthening of the above forward direction for k=n1k=n-1:
@@
Suppose there is $i_{0}\leq n$ such that $\kappa_{i_{0}}<\aleph_{i_{0}}$. Then an (n1)(n-1)-Fubini partition exists for $\vec{\kappa}$.
@@
    • If $|A|=2$, then $\operatorname{vmin}\vec{x}=\min\vec{x}$ (again, breaking ties arbitrarily).
@@
  • Third: we in fact can assume $i_{0}=n$, i.e., $\kappa_{n}<\aleph_{n}$. This is because of the general fact that an $(i_{0}-1)$-Fubini partition of $X(\kappa_{0},\ldots,\kappa_{i_{0}})^{-}$ yields an (n1)(n-1)-Fubini partition of $X(\vec{\kappa})^{-}$ by setting the remaining $Y_{i}$ equal to the empty set.
@@
So the $\langle Y_{j}\rangle$ form an (n1)(n-1)-Fubini partition.  ∎

### v2

@@ Abstract.
We compute the sheaf cohomology with constant $\mathbb{Z}_{2}$ coefficients of a concrete class of locally profinite sets of independent interest. We introduce $k$-sheer partitions to aid in constructions. It is also shown that questions of intermediate cohomology degrees can be reduced to questions about top cohomology degrees by exhibiting nontrivial top cocycles as pointwise limits of coboundaries.
@@ 1. Introduction
The paper contains three main parts, corresponding to the three main tools developed. After expositing the spaces under study in section 2, in section 3 we analyze their sheaf cohomology directly by exhibiting nontrivial cocycles via a nonconstructive argument. Here our arguments are within ZFC, but our conclusions are strengthened by cardinal arithmetic bounds. In section 4, we define $k$-sheer partitions of the spaces, which aid in constructions of cocycles and trivializers. We use these partitions to show both vanishing and nonvanishing of certain cohomology groups; we also prove a converse to one of Aoki’s main results. Lastly, in section 5 we characterize nonvanishing cohomology purely in terms of nonvanishing top cohomology. A better understanding of these top cohomology groups is thus sufficient to resolve all remaining cases. All our arguments may be construed as taking place in ZF, with choice only needed in a few places, such as in the existence of an injection $\aleph_{1}\hookrightarrow 2^{\aleph_{0}}$.
@@ 2. Definitions and overview of results
In section 4, we describe techniques for eliminating the cardinal arithmetic assumptions of Theorem A in certain situations. The main tool is a partition of the product $\prod_{i\leq n}\kappa_{i}$ into $n+1$ pieces, each of which is “thin” in one of the coordinate directions; we call such partitions sheer.
@@
Let $\vec{\kappa}$ be a length-$(n+1)$ tuple of cardinals which admits a $k$-sheer partition. Then:
@@
We determine exactly when $k$-sheer partitions exist, in the process deriving a converse to one of Aoki’s results:
@@ 3. Proof of Theorem A
    If $A\subseteq A^{\prime}$, then $\mathcal{D}_{A^{\prime}}\subseteq\mathcal{D}_{A}$ is open dense, so that the map $f\mapsto f\restriction\mathcal{D}_{A^{\prime}}$ defines a dense embedding of topological rings $\mathcal{C}_{A}\hookrightarrow\mathcal{C}_{A^{\prime}}$. In light of this and related observations, we will frequently abuse notation when dealing with functions defined on different domains; for instance, given functions $f_{A}\in\mathcal{D}_{A}$ and $f_{B}\in\mathcal{D}_{B}$, we write $f_{A}+f_{B}$ for the function defined on the intersection $\mathcal{D}_{A}\cap\mathcal{D}_{B}=\mathcal{D}_{A\cup B}$.
@@
    This is a disjoint union of $R$ and finitely many spaces of the form $\prod_{i\in A}J_{i}$, where $A\subseteq n+1$ is of size <n+1<n+1. Since $J_{i}$ is a cofinite subset of $\alpha(\kappa_{i})$ containing $\infty$, any space of this form is finite or homeomorphic to $\mathcal{C}_{\emptyset}(\vec{\kappa}\restriction_{A})$, which by the inductive hypothesis has cardinality $\kappa_{\max A}$.
    We have thus described an injection
    $\mathcal{C}_{\emptyset}(\vec{\kappa})\hookrightarrow\prod_{i\leq n}[\kappa_{i}]^{<\aleph_{0}}\times\{0,1\}\times\prod_{A\in[n+1]^{<n+1}}\kappa_{\max A},$
@@
We are now ready to prove Theorem A. We will benefit from the following notation: for a fixed tuple $\vec{\kappa}$ of length $n+1$ and fixed $k\leq n$, write
@@
That is, if $\vec{y}\in X(\kappa_{n-k+1},\ldots,\kappa_{n})$ satisfies $y_{i}=\infty$ iff $i=j$, then $f_{A\setminus\{j\}}(\vec{x},-)$ must be constant in a neighborhood of $\vec{y}$. Thus there is $\beta(\vec{x},j)\in\kappa_{j}$ such that $f_{A\setminus\{j\}}(\vec{x},\vec{y})$ is independent of $y_{j}$ whenever $y_{j}>\beta(\vec{x},j)$. Taking
@@
we find that $\sum_{j=n-k+1}^{n}f_{A\setminus\{j\}}(\vec{x},\vec{y})$ is independent of $\vec{y}$ whenever $\vec{y}>\vec{\beta}(\vec{x})$ coordinatewise.
The above describes a map $\vec{x}\mapsto\vec{\beta}(\vec{x}):\mathcal{P}\to\mathcal{Q}$; so again by our assumption the image of this map must be bounded in $\mathcal{Q}$. Taking $\vec{\beta}$ to be a common upper bound in $\mathcal{Q}$ of all the $\vec{\beta}(\vec{x})$, we find that $\sum_{j=n-k+1}^{n}f_{A\setminus\{j\}}(\vec{x},\vec{y})$ is constant for $\vec{y}>\vec{\beta}$.
@@
4. Constructions via sheer partitions
As mentioned above, our second main computational tool is the notion of a $k$-sheer partition of $X(\vec{\kappa})^{-}$. To define this, we begin with a more careful dissection of the topology of $X(\vec{\kappa})$. We make the following notation:
@@ 4. Constructions via sheer partitions
$\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}}):=\{\vec{y}\in X(\vec{\kappa}):\vec{y}\upharpoonright_{A}=\vec{x}\upharpoonright_{A}\text{ and }y_{j}\not\in\mathcal{E}_{j}\text{ for all }j\not\in A\}.$
@@
Let $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n}\rangle$ be a tuple of cardinals, and fix $1\leq k\leq n$. We call an $(n+1)$-element partition $\langle Y_{i}:i\leq n\rangle$ of $X(\vec{\kappa})^{-}$ $k$-sheer if, for any $A\in[n+1]^{\geq k+1}$, the following criteria hold:
@@
    This reformulation also makes it clearer that $F_{A}$ is disjoint from $Y_{i}$ whenever $i\not\in A$. However, it is the more verbose condition (2) which will be most useful to us.
  • It is immediate from the definition that a $k$-sheer partition is $\ell$-sheer whenever $k\leq\ell$; moreover, if $\vec{\kappa}\neq\emptyset$ then (1)(-1)-sheer partitions never exist, and any partition is $n$-sheer. Consequently, for every nonempty tuple there is a $k_{0}$ such that a $k$-sheer partition exists if and only if $k_{0}\leq k$.
The special case k=n1k=n-1 is closely related to the “almost constant” functions considered by Aoki:
Proposition 4.4.
Let $\vec{\kappa}=\langle\kappa_{0},\ldots,\kappa_{n}\rangle$. A partition of $\prod_{i\leq n}\kappa_{i}$ is (n1)(n-1)-sheer if and only if each $Y_{i}$ is “finite in ii”, in the sense that for any $\vec{x}\in\prod_{j\neq i}\kappa_{i}$, there are only finitely many $\vec{y}\in Y_{i}$ extending $\vec{x}$.
For the forward direction, suppose the $Y_{i}$ form an (n1)(n-1)-sheer partition, fix $i\leq n$, and let $\vec{x}\in\prod_{j\neq i}\kappa_{j}$. If there are no $\vec{y}$ in $Y_{i}$ extending $\vec{x}$, then we’re done, so suppose some $\vec{y}\in Y_{i}$ extends $\vec{x}$. By condition (2), there is some finite set $\mathcal{E}_{i}\subseteq\kappa_{i}$ such that $\mathcal{N}_{A}(\vec{y},\langle\mathcal{E}_{i}\rangle)\subseteq\bigcup_{j\in A}Y_{j}$; the latter is equal to the complement of $Y_{i}$ since the $Y_{j}$ partition the product. This says exactly that if $\vec{z}$ extends $\vec{x}$ and $z_{i}\not\in\mathcal{E}_{i}$, then $\vec{z}\not\in Y_{i}$; so if $\vec{z}\in Y_{i}$ extends $\vec{x}$, then $z_{i}$ lies in the finite set $\mathcal{E}_{i}$.
For the backward direction, suppose the $Y_{j}$ are each finite in $j$ and partition the product. Again let $i\leq n$ and let $A=(n+1)\setminus\{i\}$, $x\in F_{A}$. By hypothesis, there are finitely many $\vec{y}\in Y_{i}$ which agree with $\vec{x}$ on the coordinates in $A$; we let $\mathcal{E}_{i}$ be the set of their $i$-coordinates. Then $\mathcal{N}_{A}(\vec{x},\langle\mathcal{E}_{i}\rangle)\subseteq Y_{i}^{c}=\bigcup_{j\not\in A}Y_{j}$.  ∎
We defer discussion of the (non-)existence of sheer partitions to the end of the section, focusing instead on their uses.
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-sheer partition of $X(\vec{\kappa})$. Then for any $A\in[n+1]^{\geq k+1}$ and any $i\not\in A$,
@@
Fix $A\in[n+1]^{\geq k+1}$ and $i\not\in A$. For any $B\supseteq A$ we have by Remark 4.3 that $F_{B}$ and $Y_{i}$ are disjoint; that is,
@@
As a first example of this, we construct a trivializer of a $k$-cocycle using a $k$-sheer partition:
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-sheer partition of $X(\vec{\kappa})$, and suppose $\vec{f}\in\prod_{A\in[n+1]^{k+2}}\mathcal{C}_{A}(\vec{\kappa})$ is a tuple of continuous functions. For any $B\in[n+1]^{k+1}$, any $i\leq n$, and any $\vec{x}\in Y_{i}$, define
@@
If there is a $k$-sheer partition of $X(\vec{\kappa})^{-}$ then $H^{\ell}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ for all $\ell\geq k$.
@@
Figure 2. Given a 2-cocycle $\vec{f}=\langle f_{012},f_{013},f_{023},f_{123}\rangle$ and a $k$-sheer partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle g_{B}:B\in[4]^{2}\rangle$ defines a trivializer of $\vec{f}$. (For brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
We next use a $k$-sheer partition of $X(\vec{\kappa})^{-}$ to construct an embedding of the group $H^{k}(X(\vec{\lambda})^{-},\underline{\mathbb{Z}_{2}})$ into $H^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ when $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise. A priori, one might try to construct an embedding by treating large enough coordinates as infinity and mapping unrestricted values to zero, as follows:
@@
where $\vec{x}^{\infty}$ is obtained from $\vec{x}$ by setting all $x_{i}\geq\lambda_{i}$ equal to $\infty$.
@@
Figure 3 shows an example of the naive extension of a cochain from $X(\vec{\lambda})^{-}$ to $X(\vec{\kappa})^{-}$ for $\vec{\lambda}\leq\vec{\kappa}$.
@@
Thus in order to construct an embedding, we will modify the cochain $\vec{f}^{\infty}$ to a cocycle. In fact we will show that, given the existence of a $k$-sheer partition, any $k$-cochain can be modified to a $k$-cocycle without affecting its values at any limit points.
@@
Let $\langle Y_{i}:i\leq n\rangle$ be a $k$-sheer partition of $X(\vec{\kappa})^{-}$, and let $\vec{f}\in C^{k}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})$ be a $k$-cochain. Define the modification of $\vec{f}$ as follows: for any $A\in[n+1]^{k+1}$ and any $\vec{x}\in\mathcal{D}_{A}\cap Y_{i}$, define
@@
Lastly, to see that modification preserves values at limit points, let $\vec{x}\in\mathcal{D}_{A}$ be a limit point, i.e., $\vec{x}\in F_{B}$ for some $A\subseteq B\subsetneq n+1$. Since $|B|\geq k+1$, condition (2) of the $k$-sheer partition implies the existence of a neighborhood of $\vec{x}$ in which every point lies in $Y_{i}$ for some $i\in A$. Thus, in this neighborhood $f^{\operatorname{mod}}_{A}\equiv f_{A}.$  ∎
@@
Figure 4. Given a 1-cochain $\vec{f}=\langle f_{01},f_{02},f_{03},f_{12},f_{13},f_{23}\rangle$ and a $k$-sheer partition $\langle Y_{0},Y_{1},Y_{2},Y_{3}\rangle$, the above piecewise definition of $\langle f^{\operatorname{mod}}_{A}:A\in[4]^{2}\rangle$ modifies $\vec{f}$ to a cocycle without affecting its values at any limit points. (Again, for brevity the columns are labeled $Y_{i}$ rather than $\mathcal{D}_{A}\cap Y_{i}$.)
@@
Suppose $X(\vec{\kappa})^{-}$ admits a $k$-sheer partition. Then for any $\vec{\lambda}\leq\vec{\kappa}$ coordinate-wise, there is an embedding
@@
It now remains to determine when sheer partitions exist. As we saw above, the interesting cases are $0\leq k<n$. We first consider when k<n1k<n-1.
@@
Let $\vec{\kappa}$ be a weakly increasing tuple of cardinals, $0\leq k<n-1$. Then a $k$-sheer partition of $X(\vec{\kappa})^{-}$ exists if and only if $\kappa_{n-k}=\aleph_{0}$, that is, the first nk+1n-k+1 entries of $\vec{\kappa}$ are all $\aleph_{0}$.
For the forward direction, suppose $\vec{\kappa}$ admits a $k$-sheer partition. Then we simultaneously have that (i) $\vec{\kappa}$ also admits a $(k+1)$-sheer partition (Remark 4.3), and that (ii) $H^{k+1}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})=0$ (Corollary 4.7). The only way to reconcile these with Proposition 4.12 is if no $\vec{\lambda}\leq\vec{\kappa}$ has nonzero $(k+1)$-cohomology. But now consider the tuple
@@
Then $\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$ is a neighborhood of $\vec{x}$, all of whose points have the same minimum. If $\vec{x}\in Y_{i}$ then this neighborhood lies within $Y_{i}$ and if $\vec{x}\not\in Y_{i}$ it lies outside $Y_{i}$. So $Y_{i}$ is clopen in $\mathcal{D}_{A}$. A similar approach shows (2): fix $\vec{x}\in F_{A}$ and for $j\not\in A$ set $\mathcal{E}_{j}=\{0,1,\ldots,x_{i_{0}}\}$. If $\vec{y}\in\mathcal{N}_{A}(\vec{x},\vec{\mathcal{E}})$, then for every $j\not\in A$ we have $y_{j}>x_{i_{0}}=y_{i_{0}}$; so $\min\vec{y}\neq y_{j}$ and thus $\vec{y}\not\in Y_{j}$. We conclude that the minimum of $\vec{y}$ is among the $y_{i}$ for $i\in A$, i.e.,
@@
The remaining case is k=n1k=n-1. Here the backward direction of Proposition 4.13 still holds: if $\kappa_{1}=\aleph_{0}$ then the above defines an (n1)(n-1)-sheer partition, thus trivializing the top cohomology of $X(\vec{\kappa})^{-}$. On the other hand, Aoki showed that if $\kappa_{i}\geq\aleph_{i}$ for all $i$, then $H^{n}(X(\vec{\kappa})^{-},\underline{\mathbb{Z}_{2}})\neq 0$. Our converse to Aoki’s result comes from the requisite strengthening of the above forward direction for k=n1k=n-1:
@@
Suppose there is $i_{0}\leq n$ such that $\kappa_{i_{0}}<\aleph_{i_{0}}$. Then an (n1)(n-1)-sheer partition exists for $\vec{\kappa}$.
@@
    • If $|A|=2$, then $\operatorname{vmin}\vec{x}=\min\vec{x}$.
@@
  • Third: we in fact can assume $i_{0}=n$, i.e., $\kappa_{n}<\aleph_{n}$. This is because of the general fact that an $(i_{0}-1)$-sheer partition of $X(\kappa_{0},\ldots,\kappa_{i_{0}})^{-}$ yields an (n1)(n-1)-sheer partition of $X(\vec{\kappa})^{-}$ by setting the remaining $Y_{i}$ equal to the empty set.
@@
So the $\langle Y_{j}\rangle$ form an (n1)(n-1)-sheer partition.  ∎